# LongCat-Flash-Omni Technical Report

### v1

@@ 7.3.1 Base Model Evaluation
We compare the LongCat-Flash-Omni Base model (i.e., after pre-training stage 5) with other strong text-based models across the following core capabilities and corresponding benchmarks.: (1) General Tasks: MMLU (Hendrycks et al., 2021a), MMLU-Pro (Wang et al., 2024b), C-Eval (Huang et al., 2023), and CMMLU (Li et al., 2023). (2) Reasoning Tasks: GPQA (Rein et al., 2023), SuperGPQA (M-A-P Team, ByteDance., 2025), BBH (Suzgun et al., 2023), PIQA (Bisk et al., 2019), DROP (Dua et al., 2019), CLUEWSC (Xu et al., 2020), and WinoGrande (Sakaguchi et al., 2019). (3) Math Tasks: GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b). (4) Coding Tasks: MBPP+ (Liu et al., 2024f), HumanEval+ (Liu et al., 2024f), MultiPL-E (Cassano et al., 2022), and CRUXEval (Gu et al., 2024). We follow the same evaluation protocol as in (Meituan, 2025a) to ensure maximum fairness.
@@ 7.3.2 Instruct Model Evaluation
  • General Domains: MMLU (Hendrycks et al., 2021a), MMLU-Pro (Wang et al., 2024b), CEval (Huang et al., 2023), and CMMLU (Li et al., 2023).
  • Instruction Following: IFEval (Zhou et al., 2023), COLLIE (Yao et al., 2024), and Meeseeks (Wang et al., 2025b), Meeseeks evaluates models’ instruction-following capabilities in multi-turn scenarios through an iterative feedback framework that simulates realistic human-LLM interactions, enabling models to self-correct based on turn-specific failures and better reflect real-world usage patterns.
  • Mathematical Reasoning: MATH500 (Lightman et al., 2023), AIME24 (MAA, 2024), and BeyondAIME (ByteDance-Seed, 2025).
  • General Reasoning: GPQA-diamond (Rein et al., 2023), DROP (Dua et al., 2019), ZebraLogic (Lin et al., 2025), and GraphWalks (OpenAI, 2025a).
  • Coding: Humaneval+ (Liu et al., 2024f), MBPP+ (Liu et al., 2024f), and LiveCodeBench (2024.08-2025.05) (Jain et al., 2025).
@@
We compare LongCat-Flash-Omni with some representative text chat models including DeepSeek-V3.1 (DeepSeek-AI et al., 2025), Qwen3-235B-A22B (2507 version) (Yang et al., 2025), Kimi-K2 (MoonshotAI, 2025), GPT-4.1 (OpenAI, 2025b), Claude Sonnet-4 (Anthropic, 2025), Gemini2.5-Flash (Comanici et al., 2025) and LongCat-Flash(Meituan, 2025a). For closed-source models, we conduct evaluations through their official APIs. For models supporting both thinking and non-thinking modes (Qwen3-235B-A22B, Gemini-2.5-Flash, and Claude Sonnet-4), we explicitly configure these models to operate in non-thinking mode for a fair comparison. The evaluation results are presented in Table 13. the comprehensive evaluation demonstrates that LongCat-Flash-Omni maintains superior text capability, with consistently leading performance in different domains. Compared with LongCat-Flash, whose early base model serves as the foundation for LongCat-Flash-Omni, the latter not only exhibits no degradation in text capabilities but even achieves superior performance in certain domains. This highlights the effectiveness of our training strategy and underscores the potential synergy among different modalities in omni-modal model training.
@@ 7.4.1 Cross-modality Understanding Evaluation
In this section, we evaluate cross-modal understanding using publicly available benchmarks, including OmniBench (Li et al., 2024c), WorldSense (Hong et al., 2025), and DailyOmni (Zhou et al., 2025). We use an internally corrected version of OmniBench, as the publicly released version contains scoring deficiencies. To address the limited quality and coverage of existing benchmarks, we further introduce a new benchmark, UNO-Bench (Chen et al., 2025), comprising 1,880 human-crafted questions spanning 44 task types, with 98% of the questions requiring cross-modal reasoning. We constructed this benchmark by manually annotating our in-house dataset. This approach prevents data contamination and ensures the benchmark is highly representative of real-world application scenarios. In addition to conventional multiple-choice questions, the evaluation includes innovative multi-step open-ended questions, providing a more realistic and discriminative assessment of complex reasoning abilities.
@@ 9 Contributions
Fengjiao Chen Mengshen Zhu Wenlong Zhu Yuwei Jiang
Fengyu Yang Miao Cao Xi Nan Zeyang Hu
Gan Dong Minliang Lin Xi Su Zheng Zhang
Gang Huang Nuo Xu Xiaohan Zhao Zhengkun Tian *
Gang Xu Peng Shi Xiaohao Wang Zhiqing Hong
Guanglu Wan Qi Zhang Xiaoyu Li Zhixiong Zeng
Guoqiang Tan Qian Fang Xiaoyu Wang Zhuqi Mi
Guoqiao Yu Qian Wang Xiaoyu Zhao Ziran Li
Haibo Qiu Qian Yang Xin Chen Ziwen Wang
Hao Lu Quanxiu Wang Xin Pan Ziyi Zhao
Hongbo Liu Rongxiang Weng Xiusong Sun Ziyuan Zhuang
Hongyu Xiang Rongxin Guo * Xu Xiang Zizhe Zhao
Jiaheng Wu Ruoxuan Liang Xudong Xing

### v2

@@ 7.3.1 Base Model Evaluation
We compare the LongCat-Flash-Omni Base model (i.e., after pre-training stage 5) with other strong text-based models across the following core capabilities and corresponding benchmarks.: (1) General Tasks: MMLU (hendrycks2021measuringmassivemultitasklanguage), MMLU-Pro (wang2024mmluprorobustchallengingmultitask), C-Eval (huang2023ceval), and CMMLU (li2023cmmlu). (2) Reasoning Tasks: GPQA (rein2023gpqagraduatelevelgoogleproofqa), SuperGPQA (pteam2025supergpqascalingllmevaluation), BBH (BBH), PIQA (piqa), DROP (drop), CLUEWSC (clue), and WinoGrande (winogrande). (3) Math Tasks: GSM8K (cobbe2021gsm8k), MATH (mathbenchmark). (4) Coding Tasks: MBPP+ (humanevalmbppplus), HumanEval+ (humanevalmbppplus), MultiPL-E (cassano2022multiplescalableextensibleapproach), and CRUXEval (gu2024cruxevalbenchmarkcodereasoning). We follow the same evaluation protocol as in (meituan2025longcat_flash_chat) to ensure maximum fairness.
@@ 7.3.2 Instruct Model Evaluation
  • General Domains: MMLU (hendrycks2021measuringmassivemultitasklanguage), MMLU-Pro (wang2024mmluprorobustchallengingmultitask), CEval (huang2023ceval), and CMMLU (li2023cmmlu).
  • Instruction Following: IFEval (zhou2023ifeval), COLLIE (yao2024collie), and Meeseeks (meeseeks), Meeseeks evaluates models’ instruction-following capabilities in multi-turn scenarios through an iterative feedback framework that simulates realistic human-LLM interactions, enabling models to self-correct based on turn-specific failures and better reflect real-world usage patterns.
  • Mathematical Reasoning: MATH500 (math500), AIME24 (AIME24), and BeyondAIME (bytedance_seed_2025_beyondaime).
  • General Reasoning: GPQA-diamond (rein2023gpqagraduatelevelgoogleproofqa), DROP (drop), ZebraLogic (lin2025zebralogic), and GraphWalks (graphwalks).
  • Coding: Humaneval+ (humanevalmbppplus), MBPP+ (humanevalmbppplus), and LiveCodeBench (2024.08-2025.05) (jain2025livecodebench).
@@
We compare LongCat-Flash-Omni with some representative text chat models including DeepSeek-V3.1 (deepseekai2025deepseekv3technicalreport), Qwen3-235B-A22B (2507 version) (yang2025qwen3), Kimi-K2 (Kimi_K2_web_doc), GPT-4.1 (gpt4.1), Claude Sonnet-4 (claude4), Gemini2.5-Flash (comanici2025gemini) and LongCat-Flash(meituan2025longcat_flash_chat). For closed-source models, we conduct evaluations through their official APIs. For models supporting both thinking and non-thinking modes (Qwen3-235B-A22B, Gemini-2.5-Flash, and Claude Sonnet-4), we explicitly configure these models to operate in non-thinking mode for a fair comparison. The evaluation results are presented in Table 13. the comprehensive evaluation demonstrates that LongCat-Flash-Omni maintains superior text capability, with consistently leading performance in different domains. Compared with LongCat-Flash, whose early base model serves as the foundation for LongCat-Flash-Omni, the latter not only exhibits no degradation in text capabilities but even achieves superior performance in certain domains. This highlights the effectiveness of our training strategy and underscores the potential synergy among different modalities in omni-modal model training.
@@ 7.4.1 Cross-modality Understanding Evaluation
In this section, we evaluate cross-modal understanding using publicly available benchmarks, including OmniBench (omnibench), WorldSense (worldsense), and DailyOmni (zhou2025dailyomni). We use an internally corrected version of OmniBench, as the publicly released version contains scoring deficiencies. To address the limited quality and coverage of existing benchmarks, we further introduce a new benchmark, UNO-Bench (chen2025unobench), comprising 1,880 human-crafted questions spanning 44 task types, with 98% of the questions requiring cross-modal reasoning. We constructed this benchmark by manually annotating our in-house dataset. This approach prevents data contamination and ensures the benchmark is highly representative of real-world application scenarios. In addition to conventional multiple-choice questions, the evaluation includes innovative multi-step open-ended questions, providing a more realistic and discriminative assessment of complex reasoning abilities.
@@ 9 Contributions
Fengjiao Chen Mengshen Zhu Wenlong Zhu Yuqian Li
Fengyu Yang Miao Cao Xi Nan Yuwei Jiang
Gan Dong Minliang Lin Xi Su Zeyang Hu
Gang Huang Nuo Xu Xiaohan Zhao Zheng Zhang
Gang Xu Peng Shi Xiaohao Wang Zhengkun Tian *
Guanglu Wan Qi Zhang Xiaoyu Li Zhiqing Hong
Guoqiang Tan Qian Fang Xiaoyu Wang Zhixiong Zeng
Guoqiao Yu Qian Wang Xiaoyu Zhao Zhuqi Mi
Haibo Qiu Qian Yang Xin Chen Ziran Li
Hao Lu Quanxiu Wang Xin Pan Ziwen Wang
Hongbo Liu Rongxiang Weng Xiusong Sun Ziyi Zhao
Hongyu Xiang Rongxin Guo * Xu Xiang Ziyuan Zhuang
Jiaheng Wu Ruoxuan Liang Xudong Xing Zizhe Zhao